A farm has 60 cows,pigs, and horses. There are 4 times as many cows and pigs as horses. How many horses are there?
Answer by RAY100(1637)

1) c +p+h =60, where c=#cows,,p=#pigs,,h=#horses
2) 4h = c+p
substituting (2) in (1)
4h +h =60
5h=60
h=12